OpenStudy (anonymous):

Let f(x) = (11 ln x)/(8 x) what is the derivative this is the form they want me to put it in f' = (11 + A ln x)/(B x^C)

OpenStudy (anonymous):

f(x) = 11 ln x/ 8 x = (11/8) lnx/x ->f'(x) = (11/8) ( 1 - lnx)/ x^2 = (11 - 11 lnx ) / 8x^2

OpenStudy (anonymous):

use the quotient rule: \[(du (v) - dv(u) )\div(v^2)\] Where u is the numerator and v is the denominator. du is the derivative of u and dv is the derivative of v